OpenStudy (anonymous):

Tom, wingspan and Harry started out on a journey of 100 km. Tom and Harry went by an automobile at the rate of 25km/hr, while wingspan walked at the rate of 5km/hr. After a certain distance, Harry got off and walked on at 5km/hr while Tom went back to wingspan and got him to the destination of the same time that Harry arrived. Find the number of hours required for the trip.

OpenStudy (anonymous):

I am posting the following answer.Plz check whether it is right.

OpenStudy (anonymous):

Given, total distance=100km/hr auto's speed=25km/hr wingspan's speed=5km/hr Let Harry get down from the auto after covering "x" km. Remaining distance to be covered by Harry=(100-x)km Time required=[100-x]/5 Suppose Tom meets wingspan after travelling for "y" kms. Time covered=[y/25]hrs Distance covered by wingspan=[x-y]km Time required=[(x-y)/5] hr Now, y/25=(x-y)/5 Or, 5y=25x-25y Or, 25x=30y Or, 5x=6y Or, y=5x/6 Remaining distance to be covered by Tom and wingspan=(100-y)km Speed of the auto=25km/hr Now, (100-x)/5=(600-x)/150

OpenStudy (anonymous):

In the earlier post, are all the equations correct?

OpenStudy (anonymous):

yes bhaiya u r right i m sure because each step is right here nothing is wrong

OpenStudy (anonymous):

Sad thing is that I am not going the exact answer.

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

The time to complete the trip was 8 hours. Refer to the attachment for a Mathematica solution.